OpenStudy (anonymous):

HELP!! MEDAL! I don't get LINEAR FUNCTIONS, explain plz :'( CALLING ALL MATHEMATICIANS!

OpenStudy (anonymous):

PLZ HELP

OpenStudy (anonymous):

A linear function is a function in which only the first power of the variables appears.

OpenStudy (anonymous):

For example: y=5x+4 is a linear function

OpenStudy (anonymous):

I did that Yesterday it was hard but i did it

OpenStudy (anonymous):

here is an exaple of one Example 1: Simplify the following expression. (3x + 10y) + (15x + 7y) Solution: To simplify, combine like terms. (3x + 10y) + (15x + 7y) = 3x + 10y + 15x + 7y = 3x + 15x + 10y + 7y = (3 + 15)x + (10 + 7)y = 18x + 17y

OpenStudy (anonymous):

Are you in Connections

OpenStudy (anonymous):

you just add the y's together and than ,you add the x's to gether and you have your equation.

OpenStudy (anonymous):

?? im new to this sorry but i do Algabra 7th grade and yesterday i did LINEAR FUNCTIONS

OpenStudy (anonymous):

oh.. Thank You for this information and anyone that wants to import to this can.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

7th

OpenStudy (anonymous):

who gave me a medal?

OpenStudy (anonymous):

Im in 7th to

OpenStudy (anonymous):

and if you are doing this in Study island you can click lesson next to the button and it will help you o and i gave you the Medal

OpenStudy (anonymous):

you would have made it.

OpenStudy (anonymous):

if you tryed to do it yourself ik because i did it with out any help

OpenStudy (anonymous):

I gave you one too. I am doing my lesson, I am not doing it in Study Island. Thanks!

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

can i give one to angel?

OpenStudy (anonymous):

Bye and anyone who wants to import can continue too

OpenStudy (anonymous):

k

OpenStudy (anonymous):

if you need more help ask me

OpenStudy (anonymous):

No, you can only give it to one person

OpenStudy (anonymous):

i am in a program that i can ask someone if i need help.

OpenStudy (anonymous):

but it ends at 4:30

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

how long have you been here in openstudy

OpenStudy (anonymous):

it is 3:40 now

OpenStudy (anonymous):

a few hours.

OpenStudy (anonymous):

BYE!

OpenStudy (anonymous):

i have helped a few people

OpenStudy (anonymous):

cya

OpenStudy (anonymous):

Are you going to contribute?

OpenStudy (anonymous):

Are you going to contribute?